# Constraints from Entanglement Wedge Nesting for Holography at a Finite Cutoff

### v2

@@ C.1 Derivation of Thermal RT Surface
Now that we have derived the thermal RT surface in Schwarzschild coordinates we can use the expression given for the thermal RT surface along with the transformations given in in Eq. (70). Lets begin by computing the thermal surface in the left exterior which is anchored to an interval on the boundary at time $t=t_{L}=t_{\text{Bdry}}$. We have the following expressions for $u$ and $v$
@@ C.3 Null Geodesics and Lightcones for Planar BTZ Background
Using the fact that the the Lagrangian has no explicit $x$ or $t$ dependence we have the following conserved quantities
@@ C.4 Entanglement Wedges For Constant Time Intervals in BTZ Background
An important fact about the thermal RT curves given in Eq. (66), which are anchored to constant time slice intervals on the boundary, is that they can be understood in terms of the intersection of certain lightcones whose apexes live on certain points on a conformal boundary (recall these lightcones are described by the equation given in Eq. (214)). In particular, for the thermal RT surface given in Eq. (66) we can define two lightcones. One lightcone is generated by past directed null geodesics originating from from the point $(t_{0}=t_{\text{Bdry}}+a,x_{0}=0,z=0)$ given by,
@@
It should be stressed that Figure 20 should be properly understood as a the $t=t_{\text{Br}}$ slice of the setup. When we go to a different slice $t=t_{\text{Br}}$ the dotted lines which separate $R_{L,M,R}$ remain fixed, however, the solid black line representing the brane will shift away from $z=z_{Br}$ to a different position given by Eq. (63). It will be more convenient to express the the trajectory of the brane in $z$ rather than $r$ coordinates. Using the transformation $z=L^{2}/r$ the trajectory of the brane reads

### v3

@@ C.1 Derivation of Thermal RT Surface
Now that we have derived the thermal RT surface in Schwarzschild coordinates we can use the expression given for the thermal RT surface along with the transformations given in Eq. (70). Lets begin by computing the thermal surface in the left exterior which is anchored to an interval on the boundary at time $t=t_{L}=t_{\text{Bdry}}$. We have the following expressions for $u$ and $v$
@@ C.3 Null Geodesics and Lightcones for Planar BTZ Background
Using the fact that the Lagrangian has no explicit $x$ or $t$ dependence we have the following conserved quantities
@@ C.4 Entanglement Wedges For Constant Time Intervals in BTZ Background
An important fact about the thermal RT curves given in Eq. (66), which are anchored to constant time slice intervals on the boundary, is that they can be understood in terms of the intersection of certain lightcones whose apexes live on certain points on a conformal boundary (recall these lightcones are described by the equation given in Eq. (214)). In particular, for the thermal RT surface given in Eq. (66) we can define two lightcones. One lightcone is generated by past directed null geodesics originating from the point $(t_{0}=t_{\text{Bdry}}+a,x_{0}=0,z=0)$ given by,
@@
It should be stressed that Figure 20 should be properly understood as a the $t=t_{\text{Br}}$ slice of the setup. When we go to a different slice $t=t_{\text{Br}}$ the dotted lines which separate $R_{L,M,R}$ remain fixed, however, the solid black line representing the brane will shift away from $z=z_{Br}$ to a different position given by Eq. (63). It will be more convenient to express the trajectory of the brane in $z$ rather than $r$ coordinates. Using the transformation $z=L^{2}/r$ the trajectory of the brane reads